OpenStudy (anonymous):

3SR2y^3-5ySR2y+12SR8y^3 solve the radical

OpenStudy (anonymous):

go to this www.webmath.com

OpenStudy (anonymous):

i cant it wont let me....

OpenStudy (anonymous):

hey!! riya u should go to the website i gave!

OpenStudy (anonymous):

i cant!! it wont let me

OpenStudy (anonymous):

define a radical

OpenStudy (anonymous):

and i'll solve it

OpenStudy (anonymous):

what is SR is it square root?

OpenStudy (anonymous):

probably

OpenStudy (anonymous):

yes SR is squaree root

OpenStudy (anonymous):

A radical is a root sign. Radicals require an index to define what root it is. With square roots, the two is generally left off. With cube roots (index of three) a 3 is written on top of the left side of the radical.

OpenStudy (anonymous):

okay i shall attempt to solve it ;)

OpenStudy (anonymous):

btw are you going to do A levels?

OpenStudy (anonymous):

i'm doing my AS lol

OpenStudy (anonymous):

its Y^3

OpenStudy (anonymous):

and please clarify whether the numbers before the SRs are xSR or multiples

OpenStudy (anonymous):

\[{{{3}\sqrt{2y^3}} -{{5y}\sqrt{2y}}+{12\sqrt{8y^3}}}\] \[={{{3}\sqrt{2y^3}} -{{5y}\sqrt{2y}}+{12\sqrt{2^3y^3}}}\] \[={{{3}\sqrt{2y^3}} -{{5y}\sqrt{2y}}+{24\sqrt{2y^3}}}\] \[={{{27}\sqrt{2y^3}} -{{5y}\sqrt{2y}}}\] \[={{{27y}\sqrt{2y}} -{{5y}\sqrt{2y}}}\] \[={{{22y}\sqrt{2y}} }\]

OpenStudy (anonymous):

Thank you Sheg

OpenStudy (anonymous):

yes that's it

OpenStudy (anonymous):

was about to type in the answer, but oh well :P ....not

OpenStudy (anonymous):

what's that for? A-level pure maths?